OpenStudy (anonymous):

dy/dx=x^2/y given y(3)=-1 solve for y as a function of x (don't leave complex fractions in the answer) Need help desperate, dont want to fail!!!! T.T

OpenStudy (anonymous):

Is this the complete question?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

You can solve this with the separation of variables method. Arrange the equation so you have in in the form\[f(y)dy = g(x)dx\] and then integrate both sides \[\int f(y)dy = \int g(x)dx\] Then use y(3) = -1 to find the added constant from the integration.